OpenStudy (kamille):

When x<0, then \[\sqrt{x ^{2}+\frac{ 9x ^{2} }{ 16 }}\] Is equal to: \[A \frac{ 7x }{ 4}\]\[B -\frac{ 7x }{ 4 }\] \[C \frac{ 5x }{ 4 }\] \[D - \frac{ 5x }{ 4 }\] \[E \frac{ 3|x| }{ 4 }\] I know it must be simple,but olease can anyone give me any hints?

OpenStudy (kamille):

hey @bibby

OpenStudy (kamille):

@mathslover

OpenStudy (anonymous):

First take lcm and then Convert numerator as a whole square....

OpenStudy (kamille):

sorry,but what is lcm?

mathslover (mathslover):

LCM - Lowest Common Multiple.

OpenStudy (kamille):

so: \[\sqrt{\frac{ 16x ^{2}+9x ^{2}}{ 16 }}=\sqrt{\frac{ 25x ^{2} }{ 16 }}\] Like this?

OpenStudy (kamille):

and then just take out the square root? \[\frac{ 5x }{ 4}\]

OpenStudy (kamille):

or \[-\frac{ 5x }{ 4 }\]

OpenStudy (kamille):

oh it cant be negative when you take the square root?

OpenStudy (anonymous):

it has to be negative as it is given that x is less than 0...

OpenStudy (kamille):

oh,I forgot,thanks a lot,vedic:}}}

OpenStudy (anonymous):

when we take square root we get both roots for example sqrt of 4 gives 2 or (-2)...

OpenStudy (anonymous):

u r welcome...:)